# Towards Depth Foundation Model: Recent Trends in Vision-Based Depth Estimation

### v1

@@ 5 Stereo Image Depth Estimation
Evolution of model architectures.  The CNN-based models are one of the most classic and widely used technical approaches in stereo depth estimation [96, 97, 98, 99, 39, 100, 101, 102, 103, 104]. Early methods, such as PSMNet [97], used simple convolution layers to extract features, while subsequent works adopted more complex architectures like ResNet [105], DenseNet [106], and NAS search techniques [101] to extract more robust features. Since 2020, with the success of Transformers [53] in the field of language models and the advancements of Vision Transformers [87] in computer vision, the attention paradigm was proposed [107, 108, 109, 110, 111, 112, 113]. To reduce computational complexity, [108] introduces local attention or sparse attention mechanisms. Inspired by RNNs [114] and the application of the RAFT network architecture in optical flow estimation, [115, 116, 117] construct a correlation pyramid and use an RNN model to iteratively optimize the disparity map. [117] introduces learnable update strategies that dynamically adjust the update direction through attention mechanisms. [118, 119, 120, 38] and [121] utilize monocular foundation models as priors and diffusion models as inpainting tools for stereo generation, respectively, advancing stereo depth estimation through large-scale training.
Evolution of model paradigms.  The CNN-based Cost Volume paradigm [96, 97, 98, 99, 39, 100, 101, 102, 103, 104] constructs a cost volume using 2D or 3D CNNs to represent the matching cost between the left and right images, and then perform Cost Aggregation as a post-processing step to ultimately regress the disparity map. The attention paradigm [107, 108, 109, 110, 111, 112, 113, 117] models the epipolar matching problem as a sequence-to-sequence task, utilizing Self-Attention and Cross-Attention mechanisms to capture long-range dependencies between pixels. Within a single image, the self-attention mechanism models the relationships between pixels to capture contextual information, while the Cross-Attention mechanism matches corresponding pixels between the left and right images. Specifically, each pixel in the left image interacts with all pixels in the right image to compute attention weights. The iterative optimization paradigm [115, 116, 117] dynamically updates the Cost Volume or feature maps based on the current disparity map to capture more accurate matching information. It can balance performance and speed through early stopping, making it suitable for time-sensitive applications. Additionally, this paradigm exhibits strong robustness to initial errors. With significant advancements in simulation technology, generative techniques, and monocular depth estimation methods in recent years, foundation models [122, 121, 38, 41, 123, 119, 120] in the field of stereo depth estimation are beginning to emerge.
@@ 6 Multi-View Image Depth Estimation
Evolution of model architectures.  The model architectures for multi-view depth estimation have evolved from early heuristic matching to state-of-the-art Transformer and diffusion-based models. Early approaches [124, 125] typically assume that camera parameters were already known, obtained via SfM [126, 127, 128], SLAM [129, 130], or robotic arm control [131, 132], and rely on traditional dense pixel matching techniques [133] for depth estimation. This heuristic matching lays the groundwork for future developments. With the advent of deep learning, researchers have begun employing 2D CNNs [134, 135] to process the cost volume, which requires low computational cost and enables real-time scene reconstruction. At the same time, alternative approaches utilize 3D CNNs [136, 137, 138, 44, 139, 140, 141, 142, 143, 117] to achieve more accurate depth estimation. Recent trends have started to explore the use of Transformers [144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154] or diffusion models [155], which leverage global self-attention mechanisms or diffusion-based strategies to capture richer feature representations and enhance inference quality. Evolution of method paradigms.  The method paradigms for multi-view depth estimation have evolved from early PatchMatch-based matching strategies to modern token attention mechanisms. Traditional methods [124, 125] rely on the PatchMatch algorithm [133] for pixel-level matching—a simple yet influential approach that sets the stage for subsequent innovations. With the integration of deep learning, constructing a cost volume by back-projecting multi-view features using camera parameters becomes the mainstream [136, 137, 138, 134, 44], allowing for more accurate depth regression by computing feature correlations. Given the high memory consumption of cost volumes, modern approaches [139, 140, 141, 142, 143, 117] adopt a coarse-to-fine strategy: starting with a low resolution and a large number of depth hypotheses, and progressively refining the cost volume with predictions from previous stages to achieve a balance between high resolution and high accuracy. Looking ahead, emerging research is investigating the incorporation of token attention [156, 147, 149, 157, 158, 159, 160, 161, 162, 150, 151, 152, 153, 154] mechanisms in the depth estimation process, aiming to better capture both local and global contexts and offering promising new directions for multi-view depth estimation.
@@ 7 Monocular Video Depth Estimation
Video depth estimation aims to estimate per-frame depth from a given monocular video while ensuring temporal consistency across the sequence. Compared to monocular depth estimation, its primary challenge lies in maintaining consistency over time. In contrast to multi-view stereo (MVS) [146, 147] and other multi-view depth estimation methods [144, 145], it needs to handle dynamic scenes, which imposes further challenges. Since video depth estimation requires both temporally consistent and accurate depth predictions, and its technical paradigm integrates elements from both monocular and multi-view depth estimation, we consider it the ultimate problem that a depth foundation model should address. Evolution of model architectures.  With the advancement of deep learning technologies and network architectures, various network structures have been applied in recent years to improve the temporal consistency of depth estimation. With the rise and widespread adoption of RNNs in language models, in 2019, [164] proposed using the LSTM mechanism to integrate temporal information, thereby enhancing the stability of video depth estimation. Meanwhile, in image tasks, CNN architectures (including U-Net and ResNet) have shown excellent performance. In 2020, [165] introduced test-time optimization to single-frame depth estimation methods based on CNN architectures, amplifying the capabilities of traditional CNNs through optimization and alignment. Subsequently, attention mechanisms gained significant attention.  [166, 152, 151, 153] applied attention mechanisms to integrate temporal features, maintaining memory through attention mechanisms rather than LSTMs, achieving more accurate and consistent depth estimation results. Later, with the remarkable success of diffusion models in the field of image generation and the impressive results of video diffusion models, in 2024,  [167, 168] proposed using video diffusion models to enhance consistency, achieving unprecedented stability and predictive performance.
@@
Figure 9: Representative pipelines of monocular video depth estimation. The Vision Transformer (ViT) [63] paradigm processes patchified inputs through self-attention and temporal attention mechanisms to integrate spatial and temporal depth cues [170]. The Video Diffusion paradigm leverages denoising diffusion models, using concatenated depth and image features, optionally enriched with CLIP embeddings, to generate consistent video depth estimates [168]. These scalable architectures enhance generalization and enable zero-shot depth estimation across diverse datasets.
Evolution towards foundation models.  In terms of Model Size, recent work has focused on using scalable network architectures to model video depth. Buffer Anytime [171] proposed using a Temporal ViT architecture with 343 million parameters to integrate temporal information and generate a large amount of pseudo ground truth (GT) based on single-image priors. Specifically, the input images are first patchified and converted into tokens recognizable by ViT. Then, self-attention is applied to these tokens. To fuse temporal information, cross-attention is performed on the single-frame attention results, reprocessing the features to obtain tokens integrated with temporal information. Finally, the tokens are decoded and un-patchified to produce the final depth estimation results. Another technical paradigm [167, 168, 82] proposed using video diffusion models with 200 million to 1 billion parameters to aid in recovering video depth. Specifically, these methods use the input video as a condition in the video diffusion denoising process. For the noisy depth during denoising, it is concatenated with the corresponding frame’s RGB color and, optionally, with CLIP embeddings as input to the denoising UNet. In terms of Data scale, the insufficient volume of video depth data has always been a critical issue. In 2023, NVDS [169] constructed the first large-scale, diverse video stereo depth dataset (14.2K videos, 2.24M frames) by using a video stereo matching method to generate pseudo labels. In 2024, methods like [167, 168, 82] addressed the lack of video data by leveraging video priors from video generation models. Similar to NVDS [169], DepthCrafter [168] annotated 200K videos using a video stereo matching method. Depth Any Video [82] created 40K videos, totaling 6M frames, using a game engine. Video Depth Anything [170] proposed using an image teacher to provide pseudo labels to compensate for the shortage of video depth ground truth (GT) labels.
@@ 8 Applications
For multi-view 3D reconstruction task, once the depth for each view is estimated, they can be directly fused to obtain the reconstructed scene geometry, which is usually represented as a point cloud or triangle mesh. MVSNet [138] and CasMVSNet [140] adopt fusibile [172], which first converts depth map of each view into a point cloud, then filters out points with poor consistency by projecting them into other views and checking, and finally fuses the point clouds from all views to obtain the reconstruction result of the entire scene. Simplerecon [173] and Murre [155] use TSDF fusion [174], which converts depth maps into sparse truncated SDF grids, averages them across multiple views, and finally extracts the surface using marching cubes to obtain a triangle mesh. Direct fusion methods are relatively efficient but require high-quality depth maps. NeuRIS [175] and MonoSDF [176] use the relative depth predicted from a monocular depth estimation method as a supervision signal, constraining the SDF field through a specially designed depth loss, thereby enhancing the quality of the 3D reconstruction.
@@ 8.2 Novel View Synthesis
In recent years, NeRF-based [177] and 3D Gaussian Splatting-based [178] methods have made significant progress in the task of novel view synthesis. High quality depth estimations can serve as strong priors for these models, enhancing their performance and accelerating convergence. Specifically, the depth map can be utilized as sampling guidance for NeRF-based methods [176, 179, 180] and as a coarse initialization for 3D Gaussian Splatting-based methods [181, 182, 183]. Additionally, the dense depth map can also be employed to learn a dense SDF field during the training process, allowing for the alignment of the geometry of NeRF or 3D Gaussians with the SDF field. This alignment can improve the geometry of the reconstructed results and facilitate the synthesis of high-quality novel views, particularly for perspectives that are distant from the training views. Some existing works [184, 185, 186, 187, 188] depend on lidar point clouds or RGB-D images as geometric priors. However, both lidar point clouds and RGB-D images can be costly to acquire, requiring additional sensors. High-quality depth estimations can serve as a substitute for these methods, unleashing the potential of novel view synthesis techniques.
@@ 9 Future Work
Data.  For all of the discussed depth estimation tasks, including monocular image, stereo image, multi-view image, and monocular video depth estimation, the lack of accurate, large-scale, high-quality, and high-variability data is currently the main concern for constructing and training a depth foundation model. Due to the unique nature of the depth estimation task, current approaches to acquiring data usually fall into two categories: depth sensor or synthetic rendering. For depth sensors, the main approach is to utilize LiDARs or ultrasonic devices. However, the acquired ground truth depth maps are usually incomplete or noisy due to the sensitive nature of the depth sensing devices. For synthetic data generation, there exist several attempts at curating high-quality hand-crafted, large static or dynamic 3D scenes by artists. However, these data are naturally limited to a small scale due to the amount of work required. Future works should focus on either utilizing self-supervision techniques to better transfer the knowledge of vast image and video data to the task of depth estimation, or developing a better approach for simulation and generation, providing artist-quality synthetic rendering and depth pairs to boost generalization ability.
@@ 10 Conclusion
11 Declaration
Funding and acknowledgments.  This work was partially supported by NSFC (No. 62172364, No. U24B20154, No. 62402427), Zhejiang Provincial Natural Science Foundation of China (No. LR25F020003), and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.
Competing interests.  This survey offers an analysis of recent vision-based depth estimation research and its trend towards depth foundation modeo, and does not introduce new datasets or materials, nor involve any competing interests.
@@ 11 Declaration
Author contributions.  Zhen Xu and Hongyu Zhou were responsible for the overall writing of the manuscript. Sida Peng, Yiyi Liao, Yue Wang, Ruizhen Hu, Xiaowei Zhou, and Hujun Bao provided critical supervision and guidance throughout the project, shaping its framework, refining technical discussions, and ensuring clarity and coherence, providing valuable oversight and feedback during drafting and revision. The remaining co-authors supported the work by evaluating key publications and charting the evolution timeline of depth estimation architectures across the monocular, stereo, multi-view, and monocular video depth estimation tasks to deliver a thorough survey.

### v2

@@ 5 Stereo Image Depth Estimation
Evolution of model architectures.  The CNN-based models are one of the most classic and widely used technical approaches in stereo depth estimation [127, 65, 128, 129, 130, 39, 131, 67, 132, 133, 134]. Early methods, such as PSMNet [128], used simple convolution layers to extract features, while subsequent works adopted more complex architectures like ResNet [135], DenseNet [136], and NAS search techniques [67] to extract more robust features. Since 2020, with the success of Transformers [85] in the field of language models and the advancements of Vision Transformers [118] in computer vision, the attention paradigm was proposed [66, 137, 138, 139, 140, 141, 142]. To reduce computational complexity, [137] introduces local attention or sparse attention mechanisms. Inspired by RNNs [143] and the application of the RAFT network architecture in optical flow estimation, [68, 144, 145] construct a correlation pyramid and use an RNN model to iteratively optimize the disparity map. [145] introduces learnable update strategies that dynamically adjust the update direction through attention mechanisms. [146, 147, 148, 38] and [149] utilize monocular foundation models as priors and diffusion models as inpainting tools for stereo generation, respectively, advancing stereo depth estimation through large-scale training.
Evolution of model paradigms.  The CNN-based Cost Volume paradigm [65, 128, 129, 130, 39, 131, 67, 132, 133, 134] constructs a cost volume using 2D or 3D CNNs to represent the matching cost between the left and right images, and then performs Cost Aggregation as a post-processing step to ultimately regress the disparity map. The attention paradigm [66, 137, 138, 139, 140, 141, 142, 145] models the epipolar matching problem as a sequence-to-sequence task, utilizing Self-Attention and Cross-Attention mechanisms to capture long-range dependencies between pixels. Within a single image, the self-attention mechanism models the relationships between pixels to capture contextual information, while the Cross-Attention mechanism matches corresponding pixels between the left and right images. Specifically, each pixel in the left image interacts with all pixels in the right image to compute attention weights. The iterative optimization paradigm [68, 144, 145] dynamically updates the Cost Volume or feature maps based on the current disparity map to capture more accurate matching information. It can balance performance and speed through early stopping, making it suitable for time-sensitive applications. Additionally, this paradigm exhibits strong robustness to initial errors. With significant advancements in simulation technology, generative techniques, and monocular depth estimation methods in recent years, foundation models [150, 149, 38, 41, 151, 147, 148] in the field of stereo depth estimation are beginning to emerge.
@@ 6 Multi-View Image Depth Estimation
Multi-view image depth estimation (more commonly known as multi-view stereo, MVS) aims to recover the 3D structure of a scene from multiple images captured from different viewpoints, leveraging known or estimated camera poses. This task is fundamental to 3D reconstruction and underpins applications such as robotics, AR/VR, and autonomous navigation. Compared to monocular and stereo settings, the multi-view scenario provides richer geometric cues and enables more accurate and complete scene understanding, but also introduces new challenges in terms of view selection, feature aggregation, and computational efficiency. It’s worth noting that the estimated depth maps are typically already aligned with the camera parameters from multiple viewpoints; thus, we do not distinguish between relative and metric depth estimation in this section. Over the years, the field has evolved from early heuristic and patch-based matching methods, through cost volume-based deep learning approaches with 2D and 3D CNNs, to recent advances utilizing transformers and diffusion models. These developments have not only improved the robustness and accuracy of depth estimation, for example, in sparse-view cases, but also paved the way for large-scale pretraining and the emergence of foundation models with strong generalization capabilities. In this section, we review the key technical paradigms, architectural innovations, and open challenges in multi-view image depth estimation, highlighting the trajectory towards foundation models and the remaining obstacles for real-world deployment.
@@
Evolution of model architectures.  The model architectures for multi-view depth estimation have evolved from early heuristic matching to state-of-the-art Transformer and diffusion-based models. Early approaches [152, 153] typically assume that camera parameters were already known, obtained via SfM [154, 155, 156], SLAM [157, 158], or robotic arm control [159, 160], and rely on traditional dense pixel matching techniques [161] for depth estimation. This heuristic matching lays the groundwork for future developments. With the advent of deep learning, researchers have begun employing 2D CNNs [162, 163] to process the cost volume, which requires low computational cost and enables real-time scene reconstruction. At the same time, alternative approaches utilize 3D CNNs [164, 165, 72, 44, 166, 167, 168, 169, 170, 145] to achieve more accurate depth estimation. Recent trends have started to explore the use of Transformers [171, 172, 74, 173, 174, 76, 75, 175, 176, 177, 178, 179] or diffusion models [180], which leverage global self-attention mechanisms or diffusion-based strategies to capture richer feature representations and enhance inference quality.
Evolution of method paradigms.  The method paradigms for multi-view depth estimation have evolved from early PatchMatch-based matching strategies to modern token attention mechanisms. Traditional methods [152, 153] rely on the PatchMatch algorithm [161] for pixel-level matching—a simple yet influential approach that sets the stage for subsequent innovations. With the integration of deep learning, constructing a cost volume by back-projecting multi-view features using camera parameters becomes the mainstream [164, 165, 72, 162, 44], allowing for more accurate depth regression by computing feature correlations. Given the high memory consumption of cost volumes, modern approaches [166, 167, 168, 169, 170, 145] adopt a coarse-to-fine strategy: starting with a low resolution and a large number of depth hypotheses, and progressively refining the cost volume with predictions from previous stages to achieve a balance between high resolution and high accuracy. Looking ahead, emerging research is investigating the incorporation of token attention [73, 173, 76, 181, 182, 183, 184, 185, 186, 75, 175, 176, 177, 178] mechanisms in the depth estimation process, aiming to better capture both local and global contexts and offering promising new directions for multi-view depth estimation.
@@ 7 Monocular Video Depth Estimation
Video depth estimation aims to estimate per-frame depth from a given monocular video while ensuring temporal consistency across the sequence. Compared to monocular depth estimation, its primary challenge lies in maintaining consistency over time. In contrast to multi-view stereo (MVS) [74, 173] and other multi-view depth estimation methods [171, 172], it needs to handle dynamic scenes, which imposes further challenges. Since video depth estimation requires both temporally consistent and accurate depth predictions, and its technical paradigm integrates elements from both monocular and multi-view depth estimation, we consider it the ultimate problem that a depth foundation model should address.
Evolution of model architectures.  With the advancement of deep learning technologies and network architectures, various network structures have been applied in recent years to improve the temporal consistency of depth estimation. With the rise and widespread adoption of RNNs in language models, in 2019, [188] proposed using the LSTM mechanism to integrate temporal information, thereby enhancing the stability of video depth estimation. Meanwhile, in image tasks, CNN architectures (including U-Net and ResNet) have shown excellent performance. In 2020, [77] introduced test-time optimization to single-frame depth estimation methods based on CNN architectures, amplifying the capabilities of traditional CNNs through optimization and alignment. Subsequently, attention mechanisms gained significant attention.  [78, 176, 175, 177] applied attention mechanisms to integrate temporal features, maintaining memory through attention mechanisms rather than LSTMs, achieving more accurate and consistent depth estimation results. Later, with the remarkable success of diffusion models in the field of image generation and the impressive results of video diffusion models, in 2024,  [80, 81] proposed using video diffusion models to enhance consistency, achieving unprecedented stability and predictive performance.
@@
Figure 9: Representative pipelines of monocular video depth estimation. The Vision Transformer (ViT) paradigm processes patchified inputs through self-attention and temporal attention mechanisms to integrate spatial and temporal depth cues. The Video Diffusion paradigm leverages denoising diffusion models, using concatenated depth and image features, optionally enriched with CLIP embeddings, to generate consistent video depth estimates. These scalable architectures enhance generalization and enable zero-shot depth estimation across diverse datasets.
Evolution towards foundation models.  In terms of Model Size, recent work has focused on using scalable network architectures to model video depth. Buffer Anytime [82] proposed using a Temporal ViT architecture with 343 million parameters to integrate temporal information and generate a large amount of pseudo ground truth (GT) based on single-image priors. Specifically, the input images are first patchified and converted into tokens recognizable by ViT. Then, self-attention is applied to these tokens. To fuse temporal information, cross-attention is performed on the single-frame attention results, reprocessing the features to obtain tokens integrated with temporal information. Finally, the tokens are decoded and unpatched to produce the final depth estimation results. Another technical paradigm [80, 81, 59] proposed using video diffusion models with 200 million to 1 billion parameters to aid in recovering video depth. Specifically, these methods use the input video as a condition in the video diffusion denoising process. For the noisy depth during denoising, it is concatenated with the corresponding frame’s RGB color and, optionally, with CLIP embeddings as input to the denoising UNet. In terms of Data scale, the insufficient volume of video depth data has always been a critical issue [189]. In 2023, NVDS [79] constructed the first large-scale, diverse video stereo depth dataset (14.2K videos, 2.24M frames) by using a video stereo matching method to generate pseudo labels. In 2024, methods like [80, 81, 59] addressed the lack of video data by leveraging video priors from video generation models. Similar to NVDS [79], DepthCrafter [81] annotated 200K videos using a video stereo matching method. Depth Any Video [59] created 40K videos, totaling 6M frames, using a game engine. Video Depth Anything [190] proposed using an image teacher to provide pseudo labels to compensate for the shortage of video depth ground truth (GT) labels.
@@ 8 Applications
Depth foundation models are expected to see a wide range of practical applications across computer vision, robotics, and graphics. By providing accurate, generalizable, and scalable depth estimation from monocular images, stereo pairs, multi-view images, or video sequences, these models could serve as a critical component for downstream tasks that require 3D scene understanding. In this section, we highlight several representative application domains where depth foundation models would make a significant impact, including 3D reconstruction, novel view synthesis, video world modeling, and robotics and autonomous driving.
@@
For multi-view 3D reconstruction tasks, once the depth for each view is estimated, they can be directly fused to obtain the reconstructed scene geometry, which is usually represented as a point cloud or triangle mesh. MVSNet [72] and CasMVSNet [167] adopt fusibile [71], which first converts the depth map of each view into a point cloud, then filters out points with poor consistency by projecting them into other views and checking, and finally fuses the point clouds from all views to obtain the reconstruction result of the entire scene. Simplerecon [191] and Murre [180] use TSDF fusion [192], which converts depth maps into sparse truncated SDF grids, averages them across multiple views, and finally extracts the surface using marching cubes to obtain a triangle mesh. Direct fusion methods are relatively efficient but require high-quality depth maps. NeuRIS [193] and MonoSDF [194] use the relative depth predicted from a monocular depth estimation method as a supervision signal, constraining the SDF field through a specially designed depth loss, thereby enhancing the quality of the 3D reconstruction.
@@ 8.2 Novel View Synthesis
In recent years, NeRF-based [195] and 3D Gaussian Splatting-based [196] methods have made significant progress in the task of novel view synthesis. High-quality depth estimations can serve as strong priors for these models, enhancing their performance and accelerating convergence. Specifically, the depth map can be utilized as sampling guidance for NeRF-based methods [194, 197, 198] and as a coarse initialization for 3D Gaussian Splatting-based methods [199, 200, 201]. Additionally, the dense depth map can also be employed to learn a dense SDF field during the training process, allowing for the alignment of the geometry of NeRF or 3D Gaussians with the SDF field. This alignment can improve the geometry of the reconstructed results and facilitate the synthesis of high-quality novel views, particularly for perspectives that are distant from the training views. Some existing works [202, 203, 204, 205, 206] depend on lidar point clouds or RGB-D images as geometric priors. However, both lidar point clouds and RGB-D images can be costly to acquire, requiring additional sensors. High-quality depth estimations can serve as a substitute for these methods, unleashing the potential of novel view synthesis techniques.
@@ 9 Future Work
Data.  For all of the discussed depth estimation tasks, including monocular image, stereo image, multi-view image, and monocular video depth estimation, the lack of accurate, large-scale, high-quality, and high-variability data is currently the main concern for constructing and training a depth foundation model. Due to the unique nature of the depth estimation task, current approaches to acquiring data usually fall into two categories: depth sensor or synthetic rendering. For depth sensors, the main approach is to utilize LiDARs or ultrasonic devices. However, the acquired ground truth depth maps are usually incomplete or noisy due to the sensitive nature of the depth sensing devices. For synthetic data generation, there exist several attempts at curating high-quality, hand-crafted, large static or dynamic 3D scenes by artists. However, these data are naturally limited to a small scale due to the amount of work required. Future works should focus on either utilizing self-supervision techniques to better transfer the knowledge of vast image and video data to the task of depth estimation, or developing a better approach for simulation and generation, providing artist-quality synthetic rendering and depth pairs to boost generalization ability.
@@ 10 Conclusion
Availability of data and materials
Not applicable.
Author contributions
@@
Zhen Xu and Hongyu Zhou were responsible for the overall writing of the manuscript. Sida Peng, Yiyi Liao, Yue Wang, Ruizhen Hu, Xiaowei Zhou, and Hujun Bao provided critical supervision and guidance throughout the project, shaping its framework, refining technical discussions, and ensuring clarity and coherence, providing valuable oversight and feedback during drafting and revision. The remaining co-authors supported the work by evaluating key publications and charting the evolution timeline of depth estimation architectures across the monocular, stereo, multi-view, and monocular video depth estimation tasks to deliver a thorough survey.
Acknowledgements
This research was supported by Zhejiang Provincial Natural Science Foundation of China under Grant No. LD25F030001, and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.
Declaration of competing interest
This survey offers an analysis of recent vision-based depth estimation research and its trend towards depth foundation modeo, and does not introduce new datasets or materials, nor involve any competing interests.
Author biography
{biography}
[figures/authors/zhenxu.jpg]Zhen Xu is a forth-year PhD student in Computer Science at Zhejiang University, advised by Prof. Xiaowei Zhou and Sida Peng. He obtained his bachelor’s degree in Computer Science from Zhejiang University in 2022. His current research focuses on 3D/4D neural reconstruction and rendering, and volumetric videos.
{biography}
[figures/authors/hongyuzhou.jpg]Hongyu Zhou is currently pursuing his PhD at Zhejiang University, working under the guidance of Prof. Yiyi Liao at the X-Dim Lab. Prior to this, he was a master student at Zhejiang University supervised by Prof. Deng Cai and Prof. Xiaofei at the State Key Lab of CAD&CG. He received his bachelor’s degree from Tongji University in 2020.
{biography}
[figures/authors/sidapeng.png]Sida Peng is an assistant professor (ZJU-100 Young Professor) at Zhejiang University. He received his Ph.D. degree from College of Computer Science and Technology at Zhejiang University in 2023, supervised by Prof. Xiaowei Zhou and Prof. Hujun Bao, and obtained his bachelor degree in Information Engineering from Zhejiang University in 2018. He received the 2024 CCF Outstanding Doctoral Dissertation Award and was selected as the 2022 Apple Scholar in AI/ML.
{biography}
[figures/authors/haotonglin.jpg]Haotong Lin is a fourth-year PhD student in Computer Science at Zhejiang University, advised by Prof. Xiaowei Zhou and Sida Peng. He obtained his bachelor degree in Computer Science from Zhejiang University in 2021. In the summer of 2022, he had the opportunity to engage in a highly enjoyable collaboration with Noah Snavely at Cornell University.
{biography}
[figures/authors/haoyuguo.jpg]Haoyu Guo is a research scientist at Shanghai Artificial Intelligence Laboratory. He obtained his Ph.D. degree in Computer Science at Zhejiang University in 2025. His research interests include 3D vision, spatial intelligence and world model.
{biography}
[figures/authors/jiahaoshao.jpg]Jiahao Shao is a student researcher at X-D Lab advised by Yiyi Liao. He is also lucky to have collaboration with Matteo Poggi. Previously he obtained his B.Eng. degree in Automation from Zhejiang University in 2024.
{biography}
[figures/authors/peishanyang.jpg]Peishan Yang is a Master’s student of Computer Science at Zhejiang University (ZJU). His research interests focus on 3D reconstruction and scene understanding.
{biography}
[figures/authors/qinglinyang.jpg]Qinglin Yang is currently a master’s student at the School of Software Engineering, Zhejiang University, advised by Dr. Sida Peng and Dr. Xiaowei Zhou. He received his B.E. degree in the School of Remote Sensing and Information Engineering, Wuhan University in 2024. His research interests include 3D foundation models for perception, understanding and generation.
{biography}
[figures/authors/shengmiao.jpg]Sheng Miao is a PhD student at Zhejiang University. His research interests include 3D Reconstruction, Generative Model in autonomous driving.
{biography}
[figures/authors/xingyihe.png]Xingyi He is a fourth-year Ph.D. student in Computer Science at Zhejiang University, advised by Prof. Xiaowei Zhou. He received his bachelor’s degree from Huazhong University of Science and Technology in 2021. His research interests include 3D computer vision, robotics, image matching, and 3D reconstruction.
{biography}
[figures/authors/yifanwang.jpg]Yifan Wang is a MS student at Zhejiang University. His research interests include 3D vision, Dynamic Reconstruction, Neural Rendering.
{biography}
[figures/authors/yuewang.png]Yue Wang is a Professor in the Department of Control Science and Engineering at Zhejiang University. He received his Ph.D. from Zhejiang University in 2016, advised by Prof. Rong Xiong. His research focuses on persistent robot autonomy using AI, machine learning, and computer vision. He has developed various robotic systems, some of which have been applied in industry.
{biography}
[figures/authors/ruizhenhu.jpg]Ruizhen Hu is a Distinguished Professor and Deputy Director of the Visual Computing Research Center at Shenzhen University. She received her Ph.D. in Applied Mathematics from Zhejiang University in 2015, and was previously an Assistant Researcher at SIAT. Her research focuses on computer graphics and embodied AI, especially 3D modeling and agent interaction.
{biography}
[figures/authors/yiyiliao.jpg]Yiyi Liao is an assistant professor at Zhejiang University, leading the X-D Lab. She was previously a postdoctoral researcher in the Autonomous Vision Group at the University of Tübingen and the Max Planck Institute for Intelligent Systems. She received her Ph.D. from Zhejiang University in 2018 and her B.S. from Xi’an Jiaotong University in 2013. Her research interests include 3D computer vision, scene understanding, and 3D generative models.
{biography}
[figures/authors/xiaoweizhou.jpg]Xiaowei Zhou is a Professor in the College of Computer Science and the State Key Laboratory of CAD&CG at Zhejiang University. Before joining ZJU in 2017, he was a postdoctoral researcher at the GRASP Lab, University of Pennsylvania. His research focuses on 3D computer vision, including 3D reconstruction, pose estimation, motion capture, scene understanding, and their applications in mixed reality and robotics.
{biography}
[figures/authors/hujunbao.png]Hujun Bao is a Professor at the State Key Laboratory of CAD&CG and the College of Computer Science and Technology, Zhejiang University. He received his B.Sc. (1987) and Ph.D. (1993) in mathematics and applied mathematics from Zhejiang University. He leads the 3D graphics computing group, focusing on geometry computing, 3D visual computing, and real-time rendering. His contributed to systems such as VisioniX, ACTS, and 2D-to-3D video conversion.
Graphical abstract